OpenStudy (anonymous):

how do you do long division on this? x^2/(x^2-1)

OpenStudy (anonymous):

but i need an answer that looks like 1+1/2(x-1)-1/2(x+1)

OpenStudy (anonymous):

dafudge i think its partial LOL?

OpenStudy (ajprincess):

It's x^2/x^2-1 nt x/x^2-1 @.Sam

sam (.sam.):

\[\huge \frac{x^2}{x ^{2}-1}\] \[\huge \frac{x^2}{(x+1)(x-1)}\] ----------------------------------------------------------- \[\huge \frac{x^2}{(x+1)(x-1)}=\frac{A}{x+1}+\frac{B}{(x-1)}\] \[\huge x^2=A(x-1)+B(x+1)\] When x=1 B=1/2 When x=-1 A=-1/2 \[\huge \frac{x^2}{(x+1)(x-1)}=-\frac{1}{2(x+1)}+\frac{1}{2(x-1)}\]

OpenStudy (zarkon):

\[\frac{x^2}{x^2-1}=\frac{x^2-1+1}{x^2-1}=\frac{x^2-1}{x^2-1}+\frac{1}{x^2-1}\] \[=1+\frac{1}{x^2-1}=1+\frac{1}{(x+1)(x-1)}\] then break up \(\displaystyle\frac{1}{(x+1)(x-1)}\)